OpenStudy (anonymous):

determine whether the equation defines y as a function of x. 6x^2+y^2=1. can someone help me out real quick with this one

OpenStudy (anonymous):

someone help me please

OpenStudy (anonymous):

help help !!

OpenStudy (zzr0ck3r):

is y^2 = x a function?

OpenStudy (anonymous):

it says determine whether the equation defines y as a function of x for 6x^2+y^2=1

OpenStudy (zzr0ck3r):

yes I know, but the question I asked has the same answer. so is y^2 = x a function of x?

OpenStudy (anonymous):

please show work for this one

OpenStudy (zzr0ck3r):

if you solve for y you get y = +-sqrt(x) correct?

OpenStudy (anonymous):

im lost please just work this one for me and i will figure this out

OpenStudy (zzr0ck3r):

6x^2+y^2=1 first solve for y y^2 = 1-6x^2 y = +- sqrt(1-6x^2) correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

next

OpenStudy (anonymous):

so yes right?

OpenStudy (zzr0ck3r):

so what is a function? it cant have two different y values for any x, so. in your clase if you put in, lets say x = 2 you will get two different answer mainly, -sqrt(1-6x^2) and sqrt(1-6x^2) so is it a function then?

OpenStudy (anonymous):

no?

OpenStudy (zzr0ck3r):

correct, but you need to understand why, do you?

OpenStudy (anonymous):

because got 2 different answers

OpenStudy (anonymous):

right